OpenStudy (anonymous):

The cafeteria at John’s work sells two kinds of sandwiches: chicken and steak. The chicken sandwich costs $1.00 and the steak sandwich costs $1.60. Today they made $66.80 from the 56 sandwiches sold. How many of the sandwiches sold were chicken?

OpenStudy (anonymous):

x = chicken y = steak 1x + 1.6y = 66.8 x + y = 56

OpenStudy (anonymous):

What do I insert for x or y though? :p

OpenStudy (anonymous):

solve for x in the first equation: x + 1.6y = 66.8 -1.6y x = -1.6y + 66.8 Now substitute this into the other equation

OpenStudy (anonymous):

Oh i thought the solution to that was the line below it (x+y=56) which had me puzzled for a moment haha

OpenStudy (anonymous):

Thats happened to me a lot:) x + y = 56 (-1.6y + 66.8) + y = 56 And just solve for y After you solve for y put the x into the equation and solve for x.

OpenStudy (anonymous):

Sorry **put the y value into the eqaution and solve forx

OpenStudy (anonymous):

to solve that equation, do I need to use the distributive property?

OpenStudy (anonymous):

-1.6y^2 + 66.8y = 56?

OpenStudy (anonymous):

or do I just subtract 66.8 from itself and 56.

OpenStudy (anonymous):

Sorry I took a while:/ Ok, just put it as -1.6y + 66.8 + y = 56 and solve for y. It's just combining the like terms so -1.6 + 1 = ?

OpenStudy (anonymous):

-.6y + 66.8 = 56 Do what you said now and subtract the 66.8 from itself and the 56